A piece of solid lead weighing 27.6 g at a temperature of 314 °C is placed in 276 g of liquid lead at a temperature of 370. °C. After a while, the solid melts and a completely liquid sample remains. Calculate the temperature after thermal equilibrium is reached, assuming no heat loss to the surroundings.

The enthalpy of fusion of solid lead is ΔHfus = 4.77 kJ/mol at its melting point of 328 °C, and the molar heat capacities for solid and liquid lead are Csolid = 26.9 J/mol K and Cliquid = 28.6 J/mol K.

Tfinal = ______ °C
1
Expert's answer
2019-03-25T07:31:38-0400

nsol(Pb) = msol(Pb) / M(Pb) = 27.6 g / 207 g/mol = 0.1333 mol.

nliq(Pb) = mliq(Pb) / M(Pb) = 276 g / 207 g/mol = 1.3333 mol.

Tsol1(Pb) = tsol1(Pb) + 273.15 = (314 + 273.15) K = 587.15 K.

Tliq1(Pb) = tliq1(Pb) + 273.15 = (370 + 273.15) K = 643.15 K.

Tmelt(Pb) = tmelt(Pb) + 273.15 = (328 + 273.15) K = 601.15 K.

The temperature after thermal equilibrium is Tfinal(Pb) = X.

Q1 = Cliq(Pb) × nliq(Pb) × (Tliq1(Pb) –Tfinal(Pb)) = 28.6 J/mol K × 1.3333 mol × (643.15K – Tfinal(Pb)).

Q2 = Csol(Pb) × nsol(Pb) × (Tmelt(Pb) – Tsol1(Pb)) + ΔHfus(Pb) × nsol(Pb) + Cliq(Pb) × nsol(Pb) × (Tfinal(Pb) – Tmelt(Pb)) = 26.9 J/mol K × 0.1333 mol × (601.15K – 587.15K) + 4770 J/mol × 0.1333 mol + 28.6 J/mol K × 0.1333 mol × (Tfinal(Pb) – 601.15 K).

Q1 = Q2

28.6 × 1.3333 × (643.15 – X) = 26.9 × 0.1333 mol × (601.15 – 587.15) + 4770 × 0.1333 + 28.6 × 0.1333 × (X – 601.15)

24524.84 – 38.13X = 50.20 + 635.84 + 3.81X – 2291.81

41.94X = 26130.61

X = 623.05. Tfinal(Pb) = 623.05K.

tfinal(Pb) = Tfinal(Pb) – 273.15 = (623.05 – 273.15)⁰C = 349.90 ⁰C ≈ 350⁰C.